OpenStudy (anonymous):

What is the length of line AB? a. 8 b. 8.3 c. 6.6 d. 9.7 Wait as I draw the problem.

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

The line.

OpenStudy (anonymous):

This problem is most easily solved by looking at ratios of similar triangles. First, get the length of AD, That can be derived from the Pythagorean Theorem where: (AD)^2 + 4^2 = 5^2 so, once you have that, which is part of the key, then you can use similar triangles:\[\frac{ AD }{ 5 } = \frac{ 5 }{ AB }\] Now, you will have AB.

OpenStudy (anonymous):

But, how do I figure out AD?

OpenStudy (anonymous):

The similar triangles comes about through AAA, looking at triangles ADC and ACB. The two triangles share angle A and a 90-degree angle, so the last corresponding set of angles are equal. You figure out AD from that first equation I gave to you. Get 4^2 over to the right by subtracting it from each side and then take the square root.

OpenStudy (anonymous):

Rewrite (AD)^2 + 4^2 = 5^2 by subtracting 4^2 from each side. Can you do that?

OpenStudy (anonymous):

You can make this a little easier by first resolving 4^2 and 5^2. What are 4^2 and 5^2?

OpenStudy (anonymous):

16 and 25

OpenStudy (anonymous):

Good, now get 16 over to the right by subtracting 16 from each side. Then take the square root.

OpenStudy (anonymous):

It gave me the square root of 9.

OpenStudy (anonymous):

And that would equal 3.

OpenStudy (anonymous):

Good, now all you have left to do is that second equation which is the ratios.

OpenStudy (anonymous):

The answer is 8.3 !

OpenStudy (anonymous):

Thank youuu

OpenStudy (anonymous):

uw! And thx for the recognition!